The rainy season has begun in West Texas. In the last week we only opened once at night; during that night we were able to capture the mirrors taken out for the 21 point test. Mirror alignment is now routine.
The rain is also having a small impact on the contractors working in W. Texas on the remote thermal area. The large pipe vault has become a nice reflecting pool and it is clear that we will need a french drain at the low end of this vault even though the expectation is that the vault with be mostly sealed to the runoff. The concrete vault has been completed and this week the electrical and mechanical contractors are placing conduit in the vault.
This week the Distance Measuring Interferometer (DMI) arrived from France and the engineering team in Austin made sure that it will fit into the test mass mounts which will come out in the 2nd week in July. The DMI will arrive in W. Texas on Thursday so that we can get it setup in the coming week both for the test mass and for the CCAS tower.
Work on the TCS (telescope control software) continues this week. The team hopes to be generating trajectories using RA/DEC or Alt/Az by the end of the week.
This last week the two engineering teams from Austin worked both days and nights to get their work done. The day time team was able to get the mounts for the Tip-Tilt camera, VAT and DMI on the telescope. They were able to get the VAT installed and found that the center mirror is just 30 mm off from the center of the X and Y tracker encoder ranges. We still need to determine where the encoder ranges are with respect to the upper hex which brings us to the work of the other engineering team. This team worked most of the nights this past week including weekends to get the 97 point test started. They determined that to take a set of 18 laser tracker data points at each tracker position takes far longer than anticipated and so descoped to a 21 point test instead of a 91 point test. Even with this descope the 21 point test took 15 hours of work.
This week the Austin engineering teams have departed and are working on new mounts for the test mass based on lessons learned from this past week. The data from the 21 point test needs to be analyzed and we will followup any questions that analysis may bring up with further laser tracker measurements.
The mirror teams are swapping two mirrors into the array this week and later this week the cabling teams will be putting on the clamps in the Azimuth cable wrap and fixing any cabling problems that might have developed with the full range to tracker motion that we can now make.
The contractors from Veliz Construction made substantial progress this past week. They completed the first concrete pour for the pipe and electrical vault and this week they will be setting up the forms for the second pour (which will be the walls of the vault. We will also have the electrical contractor on site to complete some Phase I punch list items and review the site for the Phase II electrical work. We will also have Gaffey Crane on site to work on the dome crane box which must have a new form factor to avoid collisions with the stinger from the Y-drive.
A great deal of effort was put into fixing small problems with the tracker software and hardware last week. By the end of the week we were able to run a mock 97 point test and even a mock trajectory. There are still some small problems using the GPS timing signal which is needed for running a true trajectory but this is a significant step forward. This week we will have two sets of engineers from Austin. One group will work on the 97 point test which will map out the tracker sphere in space and tie that sphere to the mirror truss through the use of the laser tracker. This will be done at night so that significant temperature changes do not alter the results. The corrections to tie the tracker and mirror truss spheres together will be the first components of our mount model.
The second team will be installing the mounts for the Video Alignment Telescope and the DMI on the corrector. Once these items are in place and we can run trajectories we will will begin to tie the tracker sphere to the central mirror and then refine the mount model further.
The contractors did not begin the “vault” cement pour as expected this last Friday. Delays caused by the significant heat and a small delay in the rebar arrival have put that off until this coming Wednesday.
The night operations team has been able to get all of the mirror with current electronics installed into the stack. As soon as more SAMS (Segment Alignment and Maintenance System) electronics cards come in from the contractor the remaining segments will be brought into the array.
We are experiencing exceptionally hot temperatures over the weekend and these are expected to continue through the week. These high temperatures seem to be impacting some of the electronics that allow us to align and maintain alignment of the primary mirror. The mirror team will continue to improve the actuator zero points, clocking of the segments and gaps between the sensors.
We did not complete the 97 point test last week. This was largely due to teething problems with the new tracker control software. We will continue to work these out this week in preparation for another attempt at this milestone next week. Next week we will also be installing the Video Alignment Telescope and DMI mounts on the corrector test mass to prepare for the alignment of the tracker with the primary mirror. We hope to be able to do both sets of tests next week.
The contractors working on our remote thermal area are making good progress and will continue their preparation for the “vault” cement pour which should take place this Friday. The vault will protect the power and glycol lines as they pass under the service road behind the HET.